Bazooka569:

Simplify by dividing negative three over eight divided by five over nine A. negative fifteen over seventy-two. B. fifteen over seventy-two. C. negative twenty-seven over forty. D. twenty-seven over forty.

jhonyy9:

\[\left( \frac{ -3 }{ 8 } \right)\div \left( \frac{ 5 }{ 9 } \right) = \]

jhonyy9:

how you divide two fractions ?

Mercury:

old question but will respond so this can be closed to divide these two fractions, change the middle division operator รท to a multiplication sign then flip the numerator and denominator on the second fraction then multiply across